How Do you solve the following system by substitution or elimination:
y=-3/2x
y=-2x+2

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

Since the RHS of both = y, they equal each other.
(-3/2)x = -2x + 2
-3x = -4x + 4
x = 4
-------
y = 6

Answer by ewatrrr(24785)


Hi
y=-3/2x |green line
y=-2x+2 |blue line
Sets Up well to solve by substitution for y in the 2nd equation
-(3/2)x = -2x + 2
-3x = -4x + 4
x = 4
y = -6 (y = (-3/2)*4)
The lines intersect at Pt(4,-6)